### v1

@@ 2 Results
Figure 1: Overview of FastECG. a. The architecture of the pre-trained backbone model. It consists of three convolution blocks, eight self-attention blocks, and one classification block in the base backbone. It is pre-trained on a public 12-lead ECG dataset using the supervised multi-label binary cross-entropy loss. b. Random-deactivation low-rank adaptation. On the downstream datasets, the pre-trained weights $\{W_{0}^{i}\}_{i=1}^{n}$ in the backbone are updated by the proposed random-deactivation low-rank adaptation. It randomly activates or deactivates the low-rank matrices ($\{B^{i}\}_{i=1}^{n}$ and $\{A^{i}\}_{i=1}^{n}$) in each trainable layer with a given probability $p$. Note that all the pre-trained weights are frozen during model training. All the low-rank matrices are activated and merged into the pre-trained weights in the testing stage. The merge process generates an ensemble network combining all the sub-networks produced by the random deactivation operation. c. One-shot rank allocation. The ranks of the low-rank matrices are determined by the proposed one-shot rank allocation method using only one gradient backward on the labeled samples. The matrices with high importance are allocated with a higher rank than those with low importance. d. Lightweight semi-supervised learning. During the low-rank adaptation process, unlabeled samples from the downstream datasets are combined with the labeled samples to estimate the statistics in batch-normalization layers. Subsequently, only the labeled data is forwarded to the self-attention blocks for CVDs detection, and the unlabeled data is released in GPU memory. This lightweight semi-supervised pipeline improves the model performance in a computational-efficient way.
2.1 Datasets and experiment protocols
@@
In this section, we utilized a large-scale dataset to pre-train the backbone model and evaluated the performance of our FastECG on four downstream datasets. First, an openly available ECG dataset (CODE-15%) collected by the Telehealth Center of the Universidade Federal de Minas Gerais[31, 5] is used for base backbone pre-training. The CODE-15% dataset contains 15% of the ECG data from the restricted CODE-full dataset[5, 32]. Specifically, there are 345779 ECG recordings from 233770 patients in the CODE-15% dataset, alongside six CVD labels: 1st degree AV block (1dAVb), left bundle branch block (LBBB), right bundle branch block (RBBB), sinus tachycardia (ST), atrial fibrillation (AF) and sinus bradycardia (SB). Each ECG recording lasts 7-10s, and the sampling rate is 300-600 Hz. Subsequently, we use four small datasets for downstream model retraining and evaluation: the Georgia 12-lead ECG Challenge (G12EC) database[33], the Chapman-Shaoxing database[34], the Ningbo database[35], and the Physikalisch-Technische Bundesanstalt (PTB-XL) database[36]. Specifically, the G12EC database contains 10344 ECG recordings from 10,344 people, and the PTB-XL database comprises 21837 recordings from 18885 patients. The Chapman database contains 10,646 recordings from 10646 patients, and the Ningbo database encompasses 40258 recordings from 40258 patients. Only 34,905 recordings in the Ningbo database are publicly available[33]. The recordings from the four downstream databases are around 10 seconds, and the sampling rate is 500 Hz. Additionally, each database contains over 17 different CVDs, and multiple CVDs can be identified from one ECG segment simultaneously.
@@
The base backbone model is first pre-trained on the CODE-15% dataset using the ECG recordings and the corresponding multi-label ground truths. The architecture of the backbone is shown in Fig.1. AdawW optimizer[37] is used for the pre-training process with a learning rate of 1e-3 and a batch size of 1024. Then, the pre-trained backbone is retrained on the four downstream datasets using different methods under limited supervision. Taking the G12EC database as an example, the ECG recordings are split into a training set and a held-out test set in a ratio of 0.9: 0.1. Then, the training set is divided into a labeled training set and an unlabeled training set in a ratio of 0.05: 0.95. A validation set is randomly sampled from the labeled training set and accounts for 20% of it, which is used for selecting the best-performing model during training. For model comparisons, we reproduce several baseline models in semi-supervised learning: ReMixMatch[23], FixMatch[8], FlexMatch[21], SoftMatch[24], MixedTeacher[9], Adsh[38], SAW[39], and parameter-efficient methods: LoRA[25], DyLoRA[40], AdaLoRA[26], IncreLoRA[30].
@@
We comprehensively evaluate the model performance of various methods using multiple metrics and training costs. Since multiple CVDs can be detected from one recording simultaneously, we used metrics on multi-label classification. Our metrics include ranking loss, coverage, mean average precision (MAP), macro AUC, macro $G_{\beta=2}$, and macro $F_{\beta=2}$. We set the $\beta$ value to be 2 for all the corresponding experiments following the configurations provided in ref.[41]. Lower values indicate better model performance in the first two metrics (ranking loss and coverage), whereas the inverse holds for the remaining metrics. A detailed computation process for the metrics can be found in Supplementary Materials Section 1. Additionally, we report the training costs of different models. Specifically, the GPU memory footprint (Mem), the number of trainable parameters (Params), and the averaged training time for each optimization iteration (Time/iter) are presented. The higher the number of trainable parameters, the higher the parameter storage consumption. Note that the number of trainable parameters of FastECG can be adjusted by the initial rank $r$. Lower ranks indicate fewer trainable parameters. The AdamW optimizer[37] is used under a learning rate of 1e-3 and a batch size of 64 for all the compared methods. All the experiments are conducted in a single NVIDIA A6000 graphics processing unit using the Pytorch library.
2.2 Analysis of the cardiovascular diseases detection result
@@
Table 1 shows that our proposed FastECG achieved superior detection performance on four datasets with the lowest computational costs compared with the baseline models. For example, in the G12EC dataset, FastECG with $r=16$ achieves a macro $F_{\beta=2}$ of 0.551$\pm$0.017, which is 4.1% larger than the second-best model’s (FixMatch) performance. In Supplementary Materials Section 2, we present the detection performance of different models on each CVD. The results demonstrate that FastECG ranks the best in some CVDs, such as atrial fibrillation and first-degree AV block. It also achieves comparable performance to the compared methods in the remaining CVDs. Regarding computational costs, it requires 33.5% less training time than MixedTeacher, occupies 29.3% less GPU memory than Adsh, and has only 5.8% of the trainable parameters found in them. When the initial rank $r$ decreases to 4, FastECG shows a slight performance drop in four datasets, but the number of trainable parameters further decreases to 1.8% of the baseline models. This observation indicates the stability and robustness of the FastECG under extremely low parameter budgets.
@@
We further compare the proposed FastECG with the parameter-efficient methods, which are integrated with FixMatch for parameter-efficient semi-supervised learning. For example, FixMatch with low-rank adaptation (LoRA) is denoted as ’FixMatch+LoRA’. Similar to the FastECG, their budgets for the number of trainable parameters are controlled by the initial rank $r$. As illustrated in Fig. 2, we report their macro-$F_{\beta=2}$ scores, macro-$G_{\beta=2}$ scores, and Time/iter on four datasets at sufficient ($r=16$) and limited ($r=4$) budget levels. The macro-$F_{\beta=2}$ and macro-$G_{\beta=2}$ scores of the FixMatch without parameter-efficient training (Params: 9.505M) are denoted as gray dotted lines. The experiment results indicate that FastECG consistently outperforms the other methods on four datasets at different budget levels. Under a sufficient parameter budget ($r=16$), FastECG achieves a macro-$G_{\beta=2}$ score of 0.307$\pm$0.016 on the G12EC dataset, which is 2.8% higher than the FixMatch with LoRA. When the parameter budget is limited ($r=4$), FastECG still outperforms it by 1.5%. Additionally, FastECG achieves the highest training speed and the best performance with the least trainable parameters compared to other parameter-efficient semi-supervised learning frameworks. On the four datasets, FastECG requires 50% less training time compared to other methods. This phenomenon demonstrates the computational efficiency of the proposed FastECG. The Extended Data Table Computation-Efficient Semi-Supervised Learning for ECG-based Cardiovascular Diseases Detection presents detailed comparison results on more evaluation metrics, which provides supplementary evidence on the efficiency of the proposed FastECG in CVDs detection. Paired t-tests are conducted to evaluate the significance levels of the performance difference between FastECG and the aforementioned SOTA methods (Supplementary Materials Section 3 Fig. S1). Based on the calculated two-sided $p$-value, it can be observed that FastECG outperforms the baselines at a 0.05 significance level in most datasets and evaluation metrics, which indicates a significant superiority for the proposed FastECG framework.
@@
In summary, our experiment results demonstrate the robustness and computational efficiency of the FastECG in semi-supervised cardiovascular disease detection at different parameter budget levels. In other words, FastECG can enhance the detection performance of ECG-based CVDs detection models without introducing heavy computation burdens.
@@
Table 1: Performance comparisons of FastECG and semi-supervised baselines on the base backbone. The average performance on all CVDs within each dataset is shown across six seeds. The standard deviation is also reported for the evaluation metrics.
@@
Figure 2: Comparison between FastECG and parameter-efficient semi-supervised methods on the base backbone. Circles with various colors denote different models, and their size represents the number of trainable parameters. The training time for each optimization iteration (Time/iter) of different methods is also reported. The gray dotted lines represent the performance of the FixMatch baseline without parameter-efficient training (approximately 9.505M trainable parameters). The first row of the figure presents the performance of different models with sufficient parameter budgets ($r=16$), while the second row reports their performance under limited parameter budgets ($r=4$).

### v2

@@ 2.1 AI-Enabled CVDs Prediction using ECG
Figure 1: Overview of CE-SSL. a. The architecture of the pre-trained backbone model. It consists of three convolution blocks, eight self-attention blocks, and one classification block in the base backbone. It is pre-trained on a public 12-lead ECG dataset using the supervised multi-label binary cross-entropy loss. b. Random-deactivation low-rank adaptation. On the downstream datasets, the pre-trained weights $\{W_{0}^{i}\}_{i=1}^{n}$ in the backbone are updated by the proposed random-deactivation low-rank adaptation. It randomly activates or deactivates the low-rank matrices ($\{B^{i}\}_{i=1}^{n}$ and $\{A^{i}\}_{i=1}^{n}$) in each trainable layer with a given probability $p$. Note that all the pre-trained weights are frozen during model training. All the low-rank matrices are activated and merged into the pre-trained weights in the testing stage. The merge process generates an ensemble network combining all the sub-networks produced by the random deactivation operation. c. One-shot rank allocation. The ranks of the low-rank matrices are determined by the proposed one-shot rank allocation method using only one gradient backward on the labeled samples. The matrices with high importance are allocated with a higher rank than those with low importance. d. Lightweight semi-supervised learning. During the low-rank adaptation process, unlabeled samples from the downstream datasets are combined with the labeled samples to estimate the statistics in batch-normalization layers. Subsequently, only the labeled data is forwarded to the self-attention blocks for CVDs detection, and the unlabeled data is released in GPU memory. This lightweight semi-supervised pipeline improves the model performance in a computational-efficient way.
TABLE I: Backbone model specifications. $N_{conv}$ indicates the number of convolution blocks, $N_{att}$ indicates the number of self-attention blocks, and $N_{cls}$ indicates the number of classification blocks. $C$ is the number of convolution channels. Hidden size is the hidden layer dimension of the self-attention blocks. Head Num is the number of heads in multi-head self-attention. Params is the total number of parameters in the backbone.
Backbone Size $N_{conv}$ $N_{att}$ $N_{cls}$ $C$ Hidden size Head Num Params
Base 3 8 1 256 256 16 9.505M
Medium 3 12 1 512 512 16 50.494M
Large 3 12 1 768 768 16 113.490M
3 Methodology
3.1 Backbone Model Pre-Training
The base backbone model is pre-trained on a public 12-lead ECG dataset (CODE-15%[39, 5]), where 345779 ECG recordings from 233770 patients are provided. The medium and large backbones are pre-trained on a restricted dataset with 2,322,513 ECG recordings from 1,558,772 patients (CODE-full[39, 40]). The specific settings of the backbone models with different sizes are shown in Table I. Note that multiple abnormalities could be identified from one ECG recording simultaneously, which indicates that a multi-label classification model should be implemented for ECG-based CVDs detection. As shown in Fig.1, The backbone model $M(X)$ consists of three parts: (1) Convolution blocks, (2) Self-attention blocks, and (3) Classification blocks. Specifically, the convolution blocks comprise multiple convolution layers (Conv) and batch normalization layers. The Leaky-Relu function is used as the activation function and skip-connection is implemented[41]. In addition, a simple but efficient self-attention pipeline is employed in the self-attention blocks[13] and two successive fully-connected layers with sigmoid activation are used for label prediction in the classification block. A multi-label binary cross-entropy function is employed for model training, defined as,
$\begin{split}&\mathcal{L}(Y,M(X))\\ &=-\frac{1}{BC}\sum^{B}_{i=1}\sum^{C}_{c=1}(1-y_{i,c})\log(1-p_{i,c})+y_{i,c}\log p_{i,c},\end{split}$ (1)
3.2 Random-Deactivation Low-Rank Adaptation
$h=(W_{0}+\triangle W)X=(W_{0}+BA)X,$ (2)
where $W_{0},\triangle W\in\mathbb{R}^{d_{1}\times d_{2}}$, $B\in\mathbb{R}^{d_{1}\times r}$ and $A\in\mathbb{R}^{r\times d_{2}}$, and the rank $r\ll\min(d_{1},d_{2})$. The LoRA freezes the pre-trained weight $W_{0}$ during model training and only optimizes the low-rank matrices $A$ and $B$, which greatly reduces the number of trainable parameters during model training[27]. However, the incremental updates of low-rank matrices are inadequate for achieving optimal performance on downstream datasets[42, 32]. To bridge the performance gap efficiently, we propose a novel random-deactivation low-rank adaptation (RD-LoRA) method, which randomly activates or deactivates the low-rank matrices in each trainable layer with a given probability $p$. To be specific, the forward process of the proposed RD-LoRA can be defined as,
$\begin{split}\mathbb{E}_{\delta\sim B(\delta,1-p)}\left[h_{t}\right]&=\mathbb{E}_{\delta\sim B(\delta,1-p)}\left[(W_{0}+\delta BA)X_{t}\right]\\ &=(\mathbb{E}\left[W_{0}\right]+(1-p)\mathbb{E}\left[BA\right])\mathbb{E}\left[X_{t}\right]\\ &=(W_{0}+(1-p)BA)X_{t}.\end{split}$ (5)
$W=W_{0}+(1-p)BA.$ (6)
3.3 Ensemble Optimization Properties of the RD-LoRA
$\begin{split}&\mathbb{E}_{\delta\sim B(\delta,1-p)}\left[\mathcal{L}(Y,M(X))\right]\\ &=(1-p)^{n}\mathcal{L}(Y,\prod_{i=1}^{n}(W_{0}^{i}+B^{i}A^{i})X)\\ &+\sum_{j=1}^{n}\left[p(1-p)^{n-1}\mathcal{L}(Y,\prod_{i=1,i\neq j}^{n}(W_{0}^{i}+B^{i}A^{i})W_{0}^{j})X\right]\\ &+\cdots+p^{n}\mathcal{L}(Y,\prod_{i=1}^{n}W^{i}_{0}X),\end{split}$ (7)
$\begin{split}&\mathcal{L}(Y,\mathbb{E}_{\delta\sim B(\delta,1-p)}\left[M(X_{t})\right])\\ &=\mathcal{L}\left(Y,\mathbb{E}_{\delta\sim B(\delta,1-p)}\left[\prod_{i=1}^{n}(W^{i}_{0}+\delta^{i}B^{i}A^{i})X_{t}\right]\right)\\ &=\mathcal{L}(Y,\prod_{i=1}^{n}(W^{i}_{0}+(1-p)B^{i}A^{i})X_{t}).\end{split}$ (8)
$\frac{\partial^{2}\mathcal{L}(Y,\sigma(M(X)))}{\partial M(X)_{c}^{2}}=\sigma(M(X))_{c}(1-\sigma(M(X))_{c})\geq 0,$ (9)
3.4 Efficient One-Shot Rank Allocation
$\begin{split}I(W^{i})&=\frac{1}{N_{e}}\sum_{j=1}^{N_{e}}(\mathcal{L}(Y,M(X))-\mathcal{L}_{W^{i}(j)=0}(Y,M(X)))^{2}\\ &\approx\left\|\frac{\partial\mathcal{L}(Y,M(X))}{\partial W^{i}}\odot W^{i}\right\|_{2}^{2},\end{split}$ (11)
$\begin{split}&\frac{\partial\mathcal{L}(Y,M(X))}{\partial W^{i}}\propto-\frac{1}{\eta}\triangle W^{i}=\frac{1}{\eta}(B_{t}^{i}A_{t}^{i}-B_{t+1}^{i}A_{t+1}^{i})\\ &=\frac{1}{\eta}\left[B_{t}^{i}A_{t}^{i}-(B_{t}^{i}-\eta\frac{\partial\mathcal{L}(Y,M(X))}{\partial B_{t}^{i}})(A_{t}^{i}-\eta\frac{\partial\mathcal{L}(Y,M(X))}{\partial A_{t}^{i}})\right]\\ &=B_{t}^{i}\frac{\partial\mathcal{L}(Y,M(X))}{\partial A_{t}^{i}}+\frac{\partial\mathcal{L}(Y,M(X))}{\partial B_{t}^{i}}A_{n}^{i}\\ &-\eta\frac{\partial\mathcal{L}(Y,M(X))}{\partial B_{t}^{i}}\frac{\partial\mathcal{L}(Y,M(X))}{\partial A_{t}^{i}},\end{split}$ (12)
$\begin{split}\frac{\partial\mathcal{L}(Y,M(X))}{\partial W_{0}^{i}}&\propto-\frac{1}{\eta}\triangle W_{0}^{i}=\frac{\partial\mathcal{L}(Y,M(X))}{\partial B_{0}^{i}}A_{0}^{i},\\ &\frac{\partial\mathcal{L}(Y,M(X))}{\partial A_{0}^{i}}=0,B_{0}^{i}=0,\end{split}$ (13)
3.5 Lightweight Semi-Supervised Learning
Semi-supervised learning (SSL) is an efficient tool for model performance enhancement when large-scale unlabeled data is available[44, 8]. Recently, many studies utilized label guessing and consistency regularization to further improve the model performance in SSL tasks, such as FixMatch[9], FlexMatch[18] and SoftMatch[21]. However, the above two techniques require the output predictions of the weak and strong-augmented unlabeled samples, which induces extra computation costs. Consequently, traditional SSL methods usually exhibit much higher memory costs and longer training time than naive supervised models. Here, we introduce a lightweight but effective SSL method without extensive consistency training and pseudo-label guessing.
where $N_{B}$ and $N_{U}$ are the numbers of labeled and unlabeled samples in the current mini-batch, and $\gamma=\frac{N_{B}}{N_{B}+N_{U}}$. Note that $N_{B}$ equals $N_{U}$ in this study, thus $\gamma=0.5$. With only limited labeled data $x_{b}$, the estimated mean $\mu_{B}=\frac{1}{N_{B}}\sum_{i=1}^{N_{B}}x_{b}^{i}$ and variance $\sigma_{B}=\frac{1}{N_{B}}\sum_{i=1}^{N_{B}}(x_{b}^{i}-\mu_{B})^{2}$ in traditional BN are prone to be influenced by the over-fitting problem according to the law of large numbers. On the contrary, the proposed semi-supervised BN alleviates the problem by utilizing large-scale unlabeled data $x_{u}$ for parameter estimation, which improves the model performance on unseen distributions. Since the BN layers do not exist in the self-attention and classification blocks, we only forward the labeled features to them to reduce memory cost and training time. Compared with the SOTA methods in semi-supervised learning, the proposed CE-SSL discards the label guessing and the consistency regularization modules. However, the results demonstrate that it achieves similar or even better performance on four datasets. More importantly, its computation costs are much lower than those of the SOTA methods, including less memory consumption and faster training speed.
Algorithm 1 CE-SSL algorithm
1:
2:- Labeled dataset $D_{B}=\{X_{b},Y_{b}\}$ and unlabeled dataset $D_{U}=\{X_{u}\}$;
3:- Pre-trained model $M_{0}=\{W_{0}^{i}\}_{i=1}^{n}$; Initial rank $r$; The ratio of important weights $c$; The random-deactivation probability $p$; Batch sizes of the labeled samples ($N_{B}=64$) and the unlabeled samples ($N_{U}=64$).
5:One-shot rank allocation
8:for 1 to $iteration$ do
9:     sample labeled data $\{x_{b},y_{b}\}$ from $D_{B}$;
10:     sample unlabeled data $\{x_{u}\}$ from $D_{U}$;
11:     apply data augmentation to $x_{b}$ and $x_{u}$;
12:     Lightweight semi-supervised learning
14:     release the unlabeled data $x_{u}$ in the GPU memory
15:     Random-deactivation low-rank adaptation
16:     initialize $h_{0}=x_{b}$
17:     for $i=1,2,...n$ do
18:         sample $\delta_{i}$ from the Bernoulli distribution $B(\delta,1-p)$
19:         $h_{i}=(W_{0}^{i}+\delta B^{i}A^{i})h_{i-1}$
20:     end for
22:     apply an early-stop strategy to avoid overfitting;
23:end for
3.6 Signal Pre-Processing and Data Augmentation
Artifact removal and data augmentation are two factors that play important roles in model performance. Firstly, we introduce the signal pre-processing pipeline employed in the proposed framework. The ECG recordings from the CODE-15% and CODE-full databases are first resampled to a 400Hz sampling rate following the configuration of the dataset provider[5]. The sampling rate of the recordings from the four downstream databases remains unchanged. Firstly, the length of all recordings is normalized into 6144 samples by zero-padding. Subsequently, a band-pass filter (1-47Hz) is applied to remove the power-line interference and baseline drift. Then, the pre-processed signals are normalized using z-score normalization. Secondly, CutMix[45] is employed for labeled data augmentation. Since the sample generation process of CutMix requires true labels that are absent in the unlabeled data, we employed the ECGAugment[11] for unlabeled data augmentation, which generates new samples by randomly selecting a transformation to perturb the pre-processed signals. Note that only the weak-augmentation module in the ECGAugment is employed.
4 Experiments and Datasets
@@
In this section, we utilized two large-scale datasets to pre-train the backbone and evaluated the performance of our CE-SSL on four downstream datasets. First, an openly available ECG dataset (CODE-15%) collected by the Telehealth Center of the Universidade Federal de Minas Gerais[39, 5] is used for the base backbone pre-training. The CODE-15% dataset contains 15% of the ECG data from the restricted CODE-full dataset[5, 40], which are used for pre-training the medium and large backbones. Specifically, there are 345779 ECG recordings from 233770 patients in the CODE-15% dataset, alongside six CVD labels: 1st degree AV block (1dAVb), left bundle branch block (LBBB), right bundle branch block (RBBB), sinus tachycardia (ST), atrial fibrillation (AF) and sinus bradycardia (SB). Each ECG recording lasts 7-10s, and the sampling rate is 300-600 Hz. Subsequently, we use four small datasets for downstream model retraining and evaluation: the Georgia 12-lead ECG Challenge (G12EC) database[46], the Chapman-Shaoxing database[47], the Ningbo database[48], and the Physikalisch-Technische Bundesanstalt (PTB-XL) database[49]. Specifically, the G12EC database contains 10344 ECG recordings from 10,344 people, and the PTB-XL database comprises 21837 recordings from 18885 patients. The Chapman database contains 10,646 recordings from 10646 patients, and the Ningbo database encompasses 40258 recordings from 40258 patients. Only 34,905 recordings in the Ningbo database are publicly available[46]. The recordings from the four downstream databases are around 10 seconds, and the sampling rate is 500 Hz. Additionally, each database contains over 17 different CVDs, and multiple CVDs can be identified from one ECG segment simultaneously.
@@
The base backbone model is first pre-trained on the CODE-15% dataset using the ECG recordings and the corresponding multi-label ground truths. The architecture of the backbone is shown in Fig.1. AdawW optimizer[50] is used for the pre-training process with a learning rate of 1e-3 and a batch size of 1024. Then, the pre-trained backbone is retrained on the four downstream datasets using different methods under limited supervision. Taking the G12EC database as an example, the ECG recordings are split into a training set and a held-out test set in a ratio of 0.9: 0.1. Then, the training set is divided into a labeled training set and an unlabeled training set in a ratio of 0.05: 0.95. A validation set is randomly sampled from the labeled training set and accounts for 20% of it, which is used for selecting the best-performing model during training. For model comparisons, we reproduce several baseline models in semi-supervised learning: ReMixMatch[20], FixMatch[9], FlexMatch[18], SoftMatch[21], MixedTeacher[10], Adsh[51], SAW[52]. Additionally, we integrate the state-of-the-art parameter-efficient methods (LoRA[27], DyLoRA[38], AdaLoRA[28], IncreLoRA[32]) with FixMatch for comparisons.
@@
We comprehensively evaluate the model performance of various methods using multiple metrics and training costs. Since multiple CVDs can be detected from one recording simultaneously, we used metrics on multi-label classification. Our metrics include ranking loss, coverage, mean average precision (MAP), macro AUC, macro $G_{\beta=2}$ score, and macro $F_{\beta=2}$ score. We set the $\beta$ value to be 2 for all the corresponding experiments following the configurations provided in ref.[35]. Lower values indicate better model performance in the first two metrics (ranking loss and coverage), whereas the inverse holds for the remaining metrics. A detailed computation process for the metrics can be found in Supplementary Materials Section 1. Additionally, we report the training costs of different models. Specifically, the peak GPU memory footprint during model training (Mem), the number of trainable parameters (Params), and the average training time for each optimization iteration (Time/iter) are presented. The higher the number of trainable parameters, the higher the parameter storage consumption. Note that the number of trainable parameters of CE-SSL can be adjusted by the initial rank $r$. Lower ranks indicate fewer trainable parameters. The AdamW optimizer[50] is used under a learning rate of 1e-3 and a batch size of 64 for all the compared methods. All the experiments are conducted in a single NVIDIA A6000 graphics processing unit using the Pytorch library.
5 Results and Discussion
5.1 Analysis of the CVDs Detection Results
@@
Table II shows that our proposed CE-SSL achieved superior detection performance on four datasets with the lowest computational costs compared with the baseline models. For example, in the G12EC dataset, CE-SSL with $r=16$ achieves a macro $F_{\beta=2}$ of 0.551$\pm$0.017, which is 4.1% larger than the second-best model’s (FixMatch) performance. In Supplementary Materials Section 2, we present the detection performance of different models on each CVD. The results demonstrate that CE-SSL ranks the best in some CVDs, such as atrial fibrillation and first-degree AV block. It also achieves comparable performance to the compared methods in the remaining CVDs. Regarding computational costs, it requires 33.5% less training time than MixedTeacher, occupies 29.3% less GPU memory than Adsh, and has only 5.8% of the trainable parameters found in them. When the initial rank $r$ decreases to 4, CE-SSL shows a slight performance drop in four datasets, but the number of trainable parameters further decreases to 1.8% of the baseline models. This observation indicates the stability and robustness of the CE-SSL under extremely low parameter budgets.
@@
We further compare the proposed CE-SSL with the parameter-efficient methods, which are integrated with FixMatch for parameter-efficient semi-supervised learning. For example, FixMatch with low-rank adaptation (LoRA) is denoted as ’FixMatch+LoRA’. Similar to the CE-SSL, their budgets for the number of trainable parameters are controlled by the initial rank $r$. As illustrated in Fig. 2, we report their macro-$F_{\beta=2}$ scores, macro-$G_{\beta=2}$ scores, and Time/iter on four datasets at sufficient ($r=16$) and limited ($r=4$) budget levels. The macro-$F_{\beta=2}$ and macro-$G_{\beta=2}$ scores of the FixMatch without parameter-efficient training (Params: 9.505M) are denoted as gray dotted lines. The experiment results indicate that CE-SSL consistently outperforms the other methods on four datasets at different budget levels. Under a sufficient parameter budget ($r=16$), CE-SSL achieves a macro-$G_{\beta=2}$ score of 0.307$\pm$0.016 on the G12EC dataset, which is 2.8% higher than the FixMatch with LoRA. When the parameter budget is limited ($r=4$), CE-SSL still outperforms it by 1.5%. Additionally, CE-SSL achieves the highest training speed and the best performance with the least trainable parameters compared to other parameter-efficient semi-supervised learning frameworks. On the four datasets, CE-SSL requires 50% less training time compared to other methods. This phenomenon demonstrates the computational efficiency of the proposed CE-SSL. In Supplementary Materials Table S8, we present detailed comparison results on more evaluation metrics, which provide supplementary evidence on the efficiency of the proposed CE-SSL in CVDs detection. Paired t-tests are conducted to evaluate the significance levels of the performance difference between CE-SSL and the aforementioned SOTA methods (Supplementary Materials Section 3 Fig. S1). Based on the calculated two-sided $p$-value, it can be observed that CE-SSL outperforms the baselines at a 0.05 significance level in most datasets and evaluation metrics, which indicates a significant superiority for the proposed CE-SSL framework.
@@
In summary, our experiment results demonstrate the robustness and computational efficiency of the CE-SSL in semi-supervised cardiovascular disease detection at different parameter budget levels. In other words, CE-SSL can enhance the detection performance of ECG-based CVDs detection models without introducing heavy computation burdens.
@@
TABLE II: Performance comparisons of CE-SSL and semi-supervised baselines on the base backbone. The average performance on all CVDs within each dataset is shown across six seeds. The standard deviation is also reported for the evaluation metrics.
@@
Figure 2: Comparison between CE-SSL and parameter-efficient semi-supervised methods on the base backbone. Circles with various colors denote different models, and their size represents the number of trainable parameters. The training time for each optimization iteration (Time/iter) of different methods is also reported. The gray dotted lines represent the performance of the FixMatch baseline without parameter-efficient training (approximately 9.505M trainable parameters). The first row of the figure presents the performance of different models with sufficient parameter budgets ($r=16$), while the second row reports their performance under limited parameter budgets ($r=4$).